OpenStudy (anonymous):

Question: What is the function to calculate y when the quantity = x? quantity price 100(min) $60 . . x y . . 100,000(max) $40 Please help need to solve urgently. Thank you so much in advance!

Directrix (directrix):

I do not understand the question. Are you stating that 100 units of an item cost $60 and 100,000 units of the same items cost $40?

OpenStudy (dumbcow):

looks like a demand function which has an inverse relationship y = k/x + c 60 = k/100 + c and 40 = k/100,000 +c solve for k and c k = 2002 c = 39.98 y = 2002/x + 39.98

OpenStudy (anonymous):

Thank you so much! You don't know how much you helped. By the way how how did you find k and c?

OpenStudy (dumbcow):

you could also do a log regression but with only 2 points it may not be very precise

OpenStudy (dumbcow):

i put it in my calculator and got something like y = 73.3 -2.9ln(x)

OpenStudy (dumbcow):

oh i just solved the system of 2 eqautions i solved for c in 1st equation then substituted into 2nd equation to find k

OpenStudy (anonymous):

brilliant. Can't believe I've wasted a day trying to solve this. Thank you again!